OpenStudy (anonymous):

Solve: 36 2x --- = --- 27 3 A. 4 B. 2 C. 3 D. 6

OpenStudy (anonymous):

first, you cross multiply

OpenStudy (anonymous):

3(36) = 2x(27)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

108 = 54x

OpenStudy (anonymous):

divide both side by 54 to leave the x by itself

OpenStudy (anonymous):

108/54 = 2, therefore x = 2

OpenStudy (anonymous):

check: 36/27 = 2(2)/ 3

OpenStudy (anonymous):

4/3 = 4/3